OpenStudy (anonymous):

FREE MEDAL AND FAN+TESTOMONIAL For which operations is the set {0, 1} closed? Choose all answers that are correct. A. multiplication B. division C. addition D. subtraction

OpenStudy (anonymous):

@ganeshie8 @chosenmatt

OpenStudy (anonymous):

@Catlover5925

OpenStudy (anonymous):

do u know

ganeshie8 (ganeshie8):

@Marki

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@HelpOfTheGods

OpenStudy (anonymous):

@_lilb_to_smooth_4_U_

OpenStudy (anonymous):

@TheSmartOne @TheAsker2002

TheSmartOne (thesmartone):

Sorry. I have never really learned about sets...

OpenStudy (anonymous):

ok : (

OpenStudy (anonymous):

dude

OpenStudy (anonymous):

wat

OpenStudy (anonymous):

b and d lol

OpenStudy (anonymous):

thx bro

OpenStudy (anonymous):

np

OpenStudy (anonymous):

For which operations is the set {0, 1} closed? Choose all answers that are correct. A. multiplication B. division C. addition D. subtraction try them all lets see multiplication 1x0=0 in ur set 1x1=1 in ur set 0x1=0 in ur set 0x0 =1 in ur set all values apears to be in ur set thus its closed

OpenStudy (anonymous):

b&d

OpenStudy (anonymous):

lets try division 1/0 undefined 0/1=0 1/1=1 0/0 = undefined undefined seems to be not in ur set thus its not closed

OpenStudy (anonymous):

b and d are wrong

OpenStudy (anonymous):

i know its multiplacation

OpenStudy (anonymous):

so ad

OpenStudy (anonymous):

:) good

OpenStudy (anonymous):

lets see D 0-1=-1 not in ur set

OpenStudy (anonymous):

for C 1+1=2 not in ur set

OpenStudy (anonymous):

Which equations show that the set of whole numbers is closed under multiplication? Choose all answers that are correct. A. –1 • –1 = 1 B. 0 • –1 = 0 C. 0 • 2 = 0 D. 2 • 1 = 2

OpenStudy (anonymous):

ermm have any doubts ?

OpenStudy (anonymous):

yes only A !

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is this a question for real ?

OpenStudy (anonymous):

so a for question number 1

OpenStudy (anonymous):

what are the whole numbers ? 0,1,2,3,.... ? does it includes 0 ?

OpenStudy (anonymous):

since it doesnt include 0 or negative what the answer would be ?

ganeshie8 (ganeshie8):

in some parts of the world 0 is included in whole numbers

OpenStudy (anonymous):

i know thats why i was asking :O